OpenStudy (anonymous):

integral from 0 to infinity e^-4x @AccessDenied

OpenStudy (anonymous):

tell whether it converges or diverges and if it converges evalute it

myininaya (myininaya):

so you have integrated the e^(-4x) thing right?

OpenStudy (anonymous):

yes its 1/4e^-4x

myininaya (myininaya):

-1/4e^(-4x)?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok so we don't plug in infinity we say the lim as something goes to infinity of that thing

OpenStudy (anonymous):

yes. limit as b goes to infinity we'll say

myininaya (myininaya):

what i'm saying is: \[\lim_{z \rightarrow \infty}[\frac{-1}{4}e^{-4z}-\frac{-1}{4}e^{-4(0)}]\]

myininaya (myininaya):

And oh did @AccessDenied want to do this one?

myininaya (myininaya):

I got your private message and I just ran here

myininaya (myininaya):

lol

myininaya (myininaya):

without noticing

OpenStudy (anonymous):

okay how would i evaluate that

OpenStudy (anonymous):

lol i messaaged both of you ;p

myininaya (myininaya):